OpenStudy (anonymous):

Roy drives a sedan that gets mileage of 30 miles per gallon. If fuel consumption is represented by g and mileage by m, which function represents the relationship between m and g? m = f(g) = 30 × g m = f(m) = 30 × m g = f(m) = 30 + m m = f(g) = g − 30 m = f(g) = g ÷ 30

OpenStudy (anonymous):

m = f(g) = 30 x g because your mileage depends on how many gallons you have, so m = f(g). And since you get 30 miles per gallon, so you would multiply (30 x g) to get the total mileage.

OpenStudy (anonymous):

The function f(x) = 2x2 + 3x + 5, when evaluated, gives a value of 19. What is the function’s input value? 1 -1 2 -2 -3

OpenStudy (anonymous):

your input would be the x-value in the function, so set f(x) = 19 and solve for x.

OpenStudy (anonymous):

i literally dont know how to do any math it confuses me

OpenStudy (anonymous):

its ok, just do things step by step :) and just clarifying is 2x2 2 times 2, or 2 x squared or something else?

OpenStudy (anonymous):

so that answer is positive 2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

i have more questions

OpenStudy (anonymous):

nvm on that last post haha i get it now

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Eva makes a few dozen cupcakes and sells them for $8 a dozen. Which function represents the relationship between the money Eva earns, m, and the number of cupcakes she sells, c? f(c) = c × 8 f(m) = c × 8 f(c) = m × 8 f(m) = m × 8 f(m) = c × m

OpenStudy (anonymous):

the ? are x'S

OpenStudy (anonymous):

I have this math program i gotta finish by the 5th so could you just give me the answers to all my questions ? cause i got a lot of questions just section 4 and 5

OpenStudy (anonymous):

ok, just please make sure you know how to do this too..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

f(c) = c x 8

OpenStudy (anonymous):

money she makes = 8 * # of dozen cupcakes

OpenStudy (anonymous):

Roy drives a sedan that gets mileage of 30 miles per gallon. If fuel consumption is represented by g and mileage by m, which function represents the relationship between m and g? m = f(g) = 30 × g m = f(m) = 30 × m g = f(m) = 30 + m m = f(g) = g − 30 m = f(g) = g ÷ 30

OpenStudy (anonymous):

the first answers ?'s is x so is the second answer the fourth answer is - the fifth answer is division

OpenStudy (anonymous):

you already asked this

OpenStudy (anonymous):

i accidentally closed it

OpenStudy (anonymous):

Which function represents the number of seconds, s, in d days? (Hint: The number seconds in a day is 24 × 60 × 60 = 86,400.) The function that represents the number of seconds, s, in d days is f(d) = .

OpenStudy (anonymous):

oh, m = f(g) = 30 x g

OpenStudy (anonymous):

24 x 60 x 60= 86,400 and then its a blank box i have to put in an answer

OpenStudy (anonymous):

Do you have skype or something that i can just share my screen and show you them instead of typing them?

OpenStudy (anonymous):

f(d) =s = d/(24 x 60 x 60)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can i have your skype name so i can add you ?

OpenStudy (anonymous):

or do you wanna add me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

dustin.sisler1 is my skype

OpenStudy (anonymous):

k

OpenStudy (anonymous):

did you add me?

OpenStudy (anonymous):

yep